A venture capitalist, willing to invest $1,000,000, has three investments to choose from. The first investment, a software company, has a 15% chance of returning $8,000,000 profit, a 21% chance of returning $2,500,000 profit, and a 64% chance of losing the million dollars. The second company, a hardware company, has a 16% chance of returning $4,000,000 profit, a 24% chance of returning $1,500,000 profit, and a 60% chance of losing the million dollars. The third company, a biotech firm, has a 6% chance of returning $7,000,000 profit, a 33% of no profit or loss, and a 61% chance of losing the million dollars. Order the expected values from smallest to largest. first, second, third second, first, third second, third, first first, third, second third, first, second third, second, first
Answer #1

Answer :

Software :

Particulars Profit P Expected value
Software 8,000,000

15%

=0.15

=8,000,000 * 0.15

=1,200,000

2,500,000 0.21 5,25,000
-,1,000,000 0.64 -640,000
Total

= 1,200,000 + 5,25,000 - 640,000

=$ 1,085,000

Hardware :

Particulars Profit P Expected value
Software 4,000,000

16%

=0.16

=4,000,000 * 0.16

=640,000

1,500,000 0.60 900,000
-,1,000,000 0.06 -60,000
Total

= 640,000 + 900,000 - 60,000

=$ 1,480,000

Biotech :

Particulars Profit P Expected value
Software 7,000,000

61%

=0.61

=7,000,000 * 0.61

=4,270,000

0 0.33 0
-,1,000,000 0.61 -610,000
Total

= 1,200,000 + 5,25,000 - 610,000

=$ 1,115,000

Therefore,

Order of expected value from smallest to largest is :

software smallest expected value = 1,085,000

Next Biotech = 1,115,000

Largest value hardware = 1,480,000